OpenStudy (anonymous):

How do you graph one cycle of y=(sin x/2)^2

OpenStudy (campbell_st):

well the graph moves up... the centre of the curve is y = 0.5 and the max value is y = 1 and the minvalue is y = 0 now find the period of the curve for \[y = sin(bx) \] the period is found using \[\frac{2\pi}{b}\] you have b = 0.5 or 1/2 so substitute it and find the period. that is the distance you need on the x - axis for 1 cycle hope it helps